What will happen to a spray can of paint containing only the propellant at a pressure of 750torr at 30°C if it is thrown into a heap of garbage burning at 165°C?


1
Expert's answer
2022-03-03T09:40:02-0500

P1 = 750torr

T1 = 30°C + 273 = 303k

T2 = 165°C + 273 = 438k

P2 = ?

P1/T1 = P2/T2

P2 = 750×438/303

P2 = 1084 torr

Therefore the spray can of paint will explode.